Help me break down this math problem. I know the answer is 2. I need to know how they got this answer. 2/3-x/5=4/15
Answer by rfer(16322)

-x/5=4/15-2/3
-x/5=4/15-10/15
-x/5=-6/15
-x=-6/15*5/1
-x=-30/15
x=2